OpenStudy (anonymous):

how do I make bold and italicized text in the reply box on os? bbcode doesnt work..

OpenStudy (gabbyalicorn):

Some people use LaTex and others type things into the 'Equation Editor'. Like so: \[this is italicized \]

OpenStudy (anonymous):

yeah @gabbyalicorn I understand the equation thing but how do i make pure bold inline text. and what is Latex?

OpenStudy (gabbyalicorn):

Well, you can use the latex commands for bold and underline in the equation editor. underline: \underline{text} bold: \textbf{text}

OpenStudy (anonymous):

but I can't put any whitespace in between each word :(

OpenStudy (gabbyalicorn):

This will help— \(\Huge{Font}\) tt = \(\tt{tt}\) it = \(\it{it}\) sf = \(\sf{sf}\) bf = \(\bf{bf}\) rm = \(\rm{rm}\) cal = \(\cal{cal}\) scr = \(\scr{SCR}\) mathbb = \(\mathbb{MATHBB}\) boldsymbol = \(\boldsymbol{boldsymbol}\) So some of the fonts would work you will have to write the letters in capital example: \(\mathbb{SMILE}\) = \(\mathbb{SMILE}\) The fonts that have to be capital are scr and mathbb \(\tt{Good~Luck}\) = \(\tt{Good~Luck}\)

OpenStudy (gabbyalicorn):

wait..

OpenStudy (gabbyalicorn):

bam

OpenStudy (anonymous):

ohhh so the ~ character creates a whitespace

OpenStudy (anonymous):

and parentheses keep the text inline

OpenStudy (gabbyalicorn):

Yes, if the text is out of the parenthesis, the LaTex you've made will not show in the correct form. There are many thing you can do with LaTex. Just a suggestion but, you can go into the LaTex section and find how to do all kinds of cool tricks. :)

OpenStudy (anonymous):

youre the best :D